OpenStudy (anonymous):

Find the quotient. 4a3b2+2a2b2+3ab/ab

OpenStudy (anonymous):

\(\huge \frac{4a^3b^2+2a^2b^2+3ab}{ab}=\frac{ab(4a^2b+2ab+3)}{ab}\\\Large =4a^2b+2ab+3\)